# Return-Aligned Decision Transformer

### v5

@@ 1 Introduction
Figure 1: Performance-Controllable Model. A single model adjusts its performance linearly with target returns, enabling diverse performance levels. For instance, AI opponents can be tailored to match players’ skill levels.
@@
Consider implementing AI opponents using RL in video games (Yannakakis & Togelius, 2018) and educational tools (Singla et al., 2021) development. Developers implement AI opponents with varied skill levels to ensure players from beginners to experts are appropriately challenged (Shen et al., 2021; Wu et al., 2023a). Our proposed approach allows developers to interactively adjust the target return to produce a diverse set of AI opponents with different skill levels, as illustrated in Fig. 1. This significantly reduces development overhead, enhancing game quality by providing tailored experiences for each player.
Example 2 (Agent-based modeling and simulation).
@@
Figure 2: Comparison of attention scores between DT (left) and RADT (right). DT assigns minimal attention to returns, which indicates that the action-generating features may lack significant return information. In contrast, RADT focuses attention solely on returns, generating actions based on target returns. The attention score on the right is obtained from SeqRA of RADT. A detailed analysis can be found in Sec. 3.
@@
In this paper, we propose the Return-Aligned Decision Transformer (RADT), a novel architecture designed to align the actual return with the target return. Our key idea is to separate the return-to-go sequence from the state-action sequence so that the return-to-go sequence more directly influences action generation. To achieve this, we adopt two complementary design strategies: Sequence Return Aligner (SeqRA), described in Sec. 4.1, and Stepwise Return Aligner (StepRA), discussed in Sec. 7. SeqRA processes return-to-go tokens from multiple past timesteps to capture long-term dependencies. In contrast, StepRA links each state or action token directly to its corresponding return-to-go token at the same timestep, thereby capturing their stepwise relationship. By integrating these two approaches, RADT effectively leverages return-to-go information throughout the decision-making. In our experiments, RADT significantly reduces the absolute error between the actual and target returns, achieving reductions of 54.9% and 34.4% compared to DT in the MuJoCo (Todorov et al., 2012) and Atari (Bellemare et al., 2013) domains, respectively. Ablation studies demonstrate that each strategy independently contributes to improving return alignment, and their combination further enhances the model’s ability to match the actual return to the target return.
@@ 3 Return Alignment Difficulty in DT
In order to understand the behavior of DT, we analyze how DT trained on the ant-medium-replay dataset of MuJoCo responds to the return-to-go throughout an episode. Ideally, the return-to-go should be zero at the end of the episode, which means that the actual return perfectly matches the target return. Figure 3 plots the changes in return-to-go input to the model during an episode. As shown in Figure 3 (left), DT reaches values of return-to-go that fall significantly below zero for all target returns, and it obtains actual returns that differ greatly from the target returns.
We hypothesize that the problem lies in the architectural design of DT. DT attempts to generate actions conditioned on the target return by including the return-to-go tokens in the input sequence. Specifically, it aggregates tokens within the input sequence based on the attention scores in self-attention, as shown in Eq. (4). The objective is to incorporate the target return, which is embedded in the return-to-go tokens, into the features used in the prediction head. However, when the majority of the attention scores are allocated to state or action tokens, the allocation to return-to-go tokens decreases. This happens because the sum of the attention scores is one, based on Eq. (4). This reduction in allocation diminishes the information of the target return contained within the features used in the prediction head, potentially degrading the control performance of action generation by the target return.
To confirm our hypothesis, we analyze attention scores using DT trained on the ant-medium-replay dataset from MuJoCo, which has trajectories with diverse returns. Inspired by the analysis approach presented in Kim et al. (2024), we use the attention scores from the first self-attention layer. Figure 2 (left) shows the episode averages of the attention scores for the first self-attention layer of DT. These scores represent the attention given to all tokens in the sequence against the state $s_{t}$ token, which is used for predicting action $a_{t}$. We observe from Fig. 2 (left) that the attention scores are biased towards state tokens, with little allocation to return-to-go tokens. This observation suggests that the input sequence passing through self-attention is not absorbing much information about the target return. Given that DT applies the prediction head to this input sequence, it is expected that the target return would not significantly influence the predicted action.
@@ 4 Return-Aligned Decision Transformer
As previously discussed, DT struggles to align the actual return with the target return due to the under-allocation of attention scores to the return-to-go tokens. One intuitive way to solve this problem is to structure the transformer layers such that the return-to-go tokens cannot be ignored when processing state and action tokens. To realize this intuition, we introduce Return-Aligned Decision Transformer (RADT).
@@
For practical purposes, RADT processes only the last $K$ timesteps of these sequences. In the transformer block, we first apply self-attention to the state-action sequence $\tau_{sa}$ to model dependencies within $\tau_{sa}$. We then apply our SeqRA and StepRA to $\tau_{sa}$ so that it strongly depends on the return-to-go sequence $\tau_{r}$. After the transformer layers, the action $a_{t}$ is predicted from the $s_{t}$ token in the processed state-action sequence $\tau_{sa}$ by the prediction head. The model is trained using the cross-entropy or mean-squared error loss between the predicted action and the ground truth.
@@ 5.2 Baselines and Settings
5.3 Results
Figures 8 and 9 show the absolute error between the actual return and target return, plotted for each target return, for MuJoCo and Atari, respectively. We then average the absolute errors over all target returns, and present the mean and standard error of these averages across seeds in Tab. 1 for MuJoCo and Tab. 2 for Atari. In these figures and tables, the target returns, actual returns, and absolute errors are normalized with the largest target return set to 100 and the smallest set to 0. Overall, RADT significantly reduces the absolute error compared to DT, achieving 44.6% on MuJoCo and 65.5% on Atari, as calculated from Figures 10 and 12 in Appendix A.
MuJoCo Domain.
In the MuJoCo domain, as shown in Tab. 1, RADT outperforms all baseline methods across all tasks. Furthermore, as illustrated in Fig. 8, RADT consistently achieves lower absolute error than the baselines in most target returns across all tasks. The baselines exhibit low sensitivity when the target return changes and cannot consistently reduce the absolute error under various target returns. For example, StARformer achieves low absolute error for specific target returns (50 and 66.7) in the ant-medium environment in Fig. 8. However, its overall performance still results in a high absolute error, as shown in Tab 1. In contrast, RADT consistently shows lower absolute errors across a wide range of target returns. These results suggest that RADT can effectively align the actual return with the target return in environments requiring fine-grained control with dense rewards.
@@ 5.4 Behavior of RADT in Achieving Return Alignment
For all target returns, DT exhibits a nearly linear decrease in the return-to-go, eventually reaching values far below zero. This indicates that DT does not adjust its behavior in response to decreases in the return-to-go. These findings suggest that DT generates actions largely independent of return-to-go decrease. In contrast, RADT demonstrates a gradual decrease in return-to-go that converges near zero, significantly reducing the margin by which it exceeds zero compared to DT. This suggests that RADT adaptively adjusts its actions in response to the changing return-to-go.
@@ 5.5 Ablation Study
We conduct an ablation study for the two types of return aligners that constitute RADT on the medium-replay dataset in MuJoCo. The results are summarized in Tab. 3. Each value represents the mean of the absolute errors between the actual return and the target return across multiple target returns. These values are then normalized so that the average and standard error of the DT’s results are 1.0. Table 3 reports that introducing either SeqRA or StepRA individually proves to be effective. Combining both aligners results in the smallest error across all tasks, implying that SeqRA and StepRA effectively complement each other. Please refer to Appendix. A.2 for the ablation study on Atari.
@@ 5.6 Comparison on Maximizing Expected Returns
Dataset Environment DC DT RADT (ours)
medium-replay halfcheetah $\mathbf{41.3}$ $36.6$ $\mathbf{41.3}\pm 0.30$
hopper $94.2$ $82.7$ $\mathbf{95.7}\pm 0.22$
walker2d $76.6$ $66.6$ $75.9\pm 1.55$
medium-expert halfcheetah $93.0$ $86.8$ $\mathbf{93.1}\pm 0.01$
hopper $\mathbf{110.4}$ $107.6$ $\mathbf{110.4}\pm 0.38$
walker2d $109.6$ $108.1$ $\mathbf{109.7}\pm 0.16$
medium halfcheetah $\mathbf{43.0}$ $42.6$ $42.8\pm 0.09$
hopper $92.5$ $67.6$ $\mathbf{93.7}\pm 1.95$
walker2d $\mathbf{79.2}$ $74.0$ $75.6\pm 0.57$
umaze antmaze $85.0$ $\mathbf{90.7}\pm 4.35$
umaze-diverse antmaze $78.5$ $\mathbf{80.7}\pm 2.37$
In this section, we evaluate the influence of our proposed method on maximizing expected returns. We focus on two environments: MuJoCo tasks characterized by dense rewards and AntMaze tasks featuring sparse rewards. For RADT, we train three instances with different seeds, each running 100 episodes. We report the average of normalized returns in Tab. 4. The normalized returns are computed so that 100 represents the score of an expert policy, as per Fu et al. (2020). RADT outperforms DT across all tasks in MuJoCo. Compared to DC, RADT demonstrates comparable or superior performance in 8 out of 11 MuJoCo and AntMaze tasks while maintaining performance degradation within 5% for the remaining tasks.
@@ 6 Related work
Recent studies have focused on formulating offline reinforcement learning (RL) as a problem of predicting action sequences that are conditioned by goals and rewards (Chen et al., 2021; Janner et al., 2021; Emmons et al., 2022; David et al., 2023; Schmidhuber, 2019; Srivastava et al., 2019). This approach differs from the popular value-based methods (Kumar et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2022) by modeling the relationship between rewards and actions through supervised learning. Decision transformer (DT) (Chen et al., 2021) introduces the concept of desired future returns and improves performance by training the transformer architecture (Vaswani et al., 2017) as a return-conditioned policy. Based on DT, various advancements have been proposed for introducing value functions (Yamagata et al., 2023; Gao et al., 2024; Wang et al., 2024), finetuning models online (Zheng et al., 2022), adjusting the history length (Wu et al., 2023b), and improving the transformer architecture (Shang et al., 2022; Kim et al., 2024).
@@
Some efforts focus on refining the transformer architecture for offline RL. StARformer (Shang et al., 2022) introduces two transformer architectures, one aggregates information at each step, and the other aggregates information across the entire sequence. The image encoding process is improved by dividing the observation images into patches and feeding them into the transformer to enhance step information, similar to Vision Transformer (Dosovitskiy et al., 2021). Decision ConvFormer (Kim et al., 2024) replaces attention with convolution to capture the inherent local dependence pattern of MDP. While these architectures preserve the input sequence structure of the transformer, comprising returns-to-go, states, and actions, they do not directly tackle the challenge of diminishing the influence of returns-to-go on the decision-making process. In contrast, our research specifically aims to align the actual return with the target return.
@@ 7 Discussion and Conclusion
In this paper, we proposed RADT, a novel decision-making model for aligning the actual return with the target return in offline RL. RADT splits the input sequence into return-to-go and state-action sequences, and reflects return-to-go in action generation by uniquely handling the return-to-go sequence. This unique handling includes two strategies that capture long-term dependencies and stepwise relationships within the return-to-go sequence. Experimental results demonstrated that RADT has superior aligning capabilities compared to existing DT-based models. One limitation of our method is a slight increase in computational cost compared to DT. This could potentially be improved by introducing a lightweight attention mechanism, such as Flash Attention (Dao et al., 2022) into our SeqRA. We believe that RADT’s alignment capability will improve the usability of offline RL agents in a wide range of applications: creating diverse AI opponents in video games and educational tools, controlling heterogeneous agents in simulations, and efficiently generating human motions for animation or robotics.

### v6

@@ 1 Introduction
Consider implementing AI opponents using RL in video games (Yannakakis & Togelius, 2018) and educational tools (Singla et al., 2021) development. Developers implement AI opponents with varied skill levels to ensure players, from beginners to experts, are appropriately challenged (Shen et al., 2021; Wu et al., 2023a). Our proposed approach allows developers to interactively adjust the target return to produce a diverse set of AI opponents with different skill levels. This significantly reduces development overhead, enhancing game quality by providing tailored experiences for each player.
@@
Figure 1: Comparison of attention scores between DT and RADT trained on the MuJoCo ant-medium-replay dataset. The figure shows the attention scores assigned to input sequences by the first self-attention layer in DT and the first SeqRA in RADT. Each value represents the episode average of the attention scores. DT assigns minimal attention to return-to-go tokens, while RADT assigns attention exclusively to return-to-go tokens, thereby improving return alignment in action generation.
@@
In this paper, we propose the Return-Aligned Decision Transformer (RADT), a novel architecture designed to align the actual return with the target return. Our key idea is to separate the return-to-go sequence from the state-action sequence so that the return-to-go sequence more directly influences action generation. To achieve this, we adopt two complementary design strategies: Sequence Return Aligner (SeqRA), described in Sec. 4.1, and Stepwise Return Aligner (StepRA), discussed in Sec. 4.2. SeqRA processes return-to-go tokens from multiple past timesteps to capture long-term dependencies. In contrast, StepRA links each state or action token directly to its corresponding return-to-go token at the same timestep, thereby capturing their stepwise relationship. By integrating these two approaches, RADT effectively leverages return-to-go information throughout the decision-making. In our experiments, RADT significantly reduces the absolute error between the actual and target returns, achieving reductions of 54.9% and 34.4% compared to DT in the MuJoCo (Todorov et al., 2012) and Atari (Bellemare et al., 2013) domains, respectively. Ablation studies demonstrate that each strategy independently contributes to improving return alignment, and their combination further enhances the model’s ability to match the actual return to the target return.
@@ 3 Return Alignment Difficulty in DT
In order to understand the behavior of DT, we analyze how DT trained on the ant-medium-replay dataset of MuJoCo responds to the return-to-go throughout an episode. Ideally, the return-to-go should be zero at the end of the episode, which means that the actual return perfectly matches the target return. Figure 2 plots the changes in return-to-go input to the model during an episode. As shown in Fig. 2 (left), DT reaches values of return-to-go that fall significantly below zero for all target returns, and it obtains actual returns that differ greatly from the target returns.
We hypothesize that the problem lies in the architectural design of DT. Specifically, DT conditions action predictions on target returns through return-to-go tokens within the input sequence. These tokens compete for attention allocation in DT’s self-attention mechanism, as shown in Eq. (4). When attention is predominantly given to state or action tokens, less attention is allocated to return-to-go tokens, weakening the influence of target returns on action prediction.
This hypothesis is supported by analyzing attention scores from DT’s first self-attention layer, as shown in Fig. 1 (left). We observe that attention is strongly biased toward state tokens, with minimal attention to return-to-go tokens. This tendency could arise because the training data (Fu et al., 2020) includes samples generated by Markov policies that do not take return as input, causing DT to disregard return.
@@ 4 Return-Aligned Decision Transformer
As previously discussed, DT struggles to align the actual return with the target return due to the under-allocation of attention scores to the return-to-go tokens. One intuitive way to solve this problem is to structure the transformer blocks such that the return-to-go tokens cannot be ignored when processing state and action tokens. To realize this intuition, we introduce Return-Aligned Decision Transformer (RADT).
@@
For practical purposes, RADT processes only the last $K$ timesteps of these sequences. In the transformer block, we first apply self-attention to the state-action sequence $\tau_{sa}$ to model dependencies within $\tau_{sa}$. We then apply our SeqRA and StepRA to $\tau_{sa}$ so that it strongly depends on the return-to-go sequence $\tau_{r}$. After the transformer blocks, the action $a_{t}$ is predicted from the $s_{t}$ token in the processed state-action sequence $\tau_{sa}$ by the prediction head. The model is trained using the cross-entropy or mean-squared error loss between the predicted action and the ground truth.
@@ 5.2 Baselines and Settings
5.3 Results
Figures 7 and 8 show the absolute error between the actual return and target return, plotted for each target return, for MuJoCo and Atari, respectively. We then average the absolute errors over all target returns, and present the mean and standard error of these averages across seeds in Tab. 1 for MuJoCo and Tab. 2 for Atari. In these figures and tables, the target returns, actual returns, and absolute errors are normalized with the largest target return set to 100 and the smallest set to 0. Overall, RADT significantly reduces the absolute error compared to DT, achieving 44.6% on MuJoCo and 65.5% on Atari, as calculated from Figures 9 and 11 in Appendix A.
MuJoCo Domain.
@@ 5.3 Results
In the MuJoCo domain, as shown in Tab. 1, RADT outperforms all baseline methods across all tasks. Furthermore, as illustrated in Fig. 7, RADT consistently achieves lower absolute error than the baselines in most target returns across all tasks. The baselines exhibit low sensitivity to changes in target returns and tend to bias their actual returns toward specific values. For example, significant errors are observed for baseline methods at lower target returns in the ant-medium-expert and walker2d-medium tasks. This is because the ant-medium-expert and walker2d-medium datasets contain few trajectories with low target returns. Additionally, StARformer achieves low absolute error for specific target returns (50 and 66.7) in the ant-medium environment in Fig. 7. However, its overall performance still results in a high absolute error, as shown in Tab 1. In contrast, RADT consistently shows lower absolute errors across a wide range of target returns. These results suggest that RADT can effectively align the actual return with the target return in environments requiring fine-grained control with dense rewards.
We observe that differences in the environment affect return alignment. From Fig. 10 in Appendix A.2, we observe that the actual returns are approximately 75 regardless of the target return across all methods. This is likely because the returns of the trajectories in halfcheetah-medium are biased around 75, causing the model to overfit to it. Therefore, substantial bias in the dataset’s return distribution may reduce the effectiveness of return alignment across most methods.
@@ 5.4 Behavior of RADT in Achieving Return Alignment
For all target returns, DT exhibits a nearly linear decrease in the return-to-go, eventually reaching values far below zero. This indicates that DT does not adjust its behavior in response to decreases in the return-to-go. These findings suggest that DT generates actions largely independent of the return-to-go decrease. In contrast, RADT demonstrates a gradual decrease in return-to-go that converges near zero, significantly reducing the margin by which it exceeds zero compared to DT. This suggests that RADT adaptively adjusts its actions in response to the changing return-to-go.
@@ 5.5 Ablation Study
We conduct ablation studies on the two types of return aligners comprising RADT using the medium-replay dataset in MuJoCo and the Atari dataset. The results for the medium-replay dataset in MuJoCo are summarized in Tab. 3, and those for the Atari dataset are presented in Tab. 4. Each value represents the mean of the absolute errors between the actual return and the target return across multiple target returns. These values are then normalized so that the average and standard error of the DT’s results are 1.0. The results from both Tab. 3 and Tab. 4 demonstrate that individually introducing either SeqRA or StepRA is effective. Combining both aligners results in the smallest error across all tasks, implying that SeqRA and StepRA effectively complement each other. In the MuJoCo domain, RADT w/o SeqRA outperforms RADT w/o StepRA, indicating that StepRA is particularly advantageous for continuous control tasks with dense rewards in MuJoCo. Conversely, in the Atari domain, RADT w/o StepRA performs equally or better compared to RADT w/o SeqRA, suggesting that SeqRA is beneficial for long-term credit assignment required in the Atari domain.
@@
Approach Breakout Pong Qbert Seaquest
w/o SeqRA $0.61\pm 0.60$ $0.91\pm 11.13$ $0.24\pm 0.02$ $0.95\pm 0.72$
w/o StepRA $0.57\pm 0.20$ $0.59\pm 7.16$ $0.24\pm 0.04$ $0.81\pm 1.11$
DT $1.00\pm 1.00$ $1.00\pm 1.00$ $1.00\pm 1.00$ $1.00\pm 1.00$
@@ 5.6 Comparison on Maximizing Expected Returns
Dataset Environment DC DT RADT (ours) Max Return of Dataset
medium-replay halfcheetah $\mathbf{41.3}$ $36.6$ $\mathbf{41.3}\pm 0.30$ $42.4$
hopper $94.2$ $82.7$ $\mathbf{95.7}\pm 0.22$ $98.7$
walker2d $76.6$ $66.6$ $75.9\pm 1.55$ $90.0$
medium-expert halfcheetah $93.0$ $86.8$ $\mathbf{93.1}\pm 0.01$ $92.9$
hopper $\mathbf{110.4}$ $107.6$ $\mathbf{110.4}\pm 0.38$ $116.1$
walker2d $109.6$ $108.1$ $\mathbf{109.7}\pm 0.16$ $109.1$
[…]
Dataset DC RADT (ours) Max Return of Dataset
umaze $85.0$ $\mathbf{90.7}\pm 4.35$ $100.0$
umaze-diverse $78.5$ $\mathbf{80.7}\pm 2.37$ $100.0$
[…]
Environment DC StARformer DT RADT (ours) Max Return of Dataset
Breakout $352.7$ $\mathbf{436.1}$ $267.5$ $302.6\pm 22.6$ $371.4$
Pong $106.5$ $\mathbf{110.8}$ $106.1$ $107.0\pm 2.1$ $116.7$
Qbert $\mathbf{67.0}$ $51.2$ $15.4$ $12.8\pm 3.5$ $5.8$
Seaquest $\mathbf{2.7}$ $1.7$ $2.5$ $1.2\pm 0.1$ $0.6$
In this section, we evaluate the influence of our proposed method on maximizing expected returns. We use three environments: MuJoCo tasks characterized by dense rewards, AntMaze tasks featuring sparse rewards, and Atari tasks requiring long-term credit assignment. For RADT, we train three instances with different seeds. Each instance is evaluated over 100 episodes for MuJoCo and AntMaze, and 10 episodes for Atari. The resulting average normalized returns are summarized in Tab. 5 for MuJoCo and AntMaze, and in Tab. 7 for Atari. The normalized returns are computed so that 100 represents the score of an expert policy, as per Fu et al. (2020) for MuJoCo and AntMaze, and Hafner et al. (2021) for Atari. Additionally, these tables include the highest episode returns in each dataset.
From Tab. 5 and Tab. 6, RADT achieves performance close to the maximum returns of the datasets in MuJoCo and AntMaze, and its performance is comparable to or exceeds that of DT and DC. Similarly, from Tab. 7, RADT reaches returns near the maximum values of the datasets in Breakout and Pong. In Qbert and Seaquest, although RADT also successfully achieves returns well above the maximum returns of the datasets, it falls short of the baselines specializing in maximizing returns, despite outperforming all approaches in return alignment as shown in Tab 2. RADT effectively reproduces near-maximum dataset performance, but faces challenges in scenarios requiring extrapolation to substantially higher target returns beyond the training distribution. We further discuss this limitation in Sec. 7.
@@ 6 Related work
Recent studies have focused on formulating offline reinforcement learning (RL) as a problem of predicting action sequences that are conditioned by goals and rewards (Chen et al., 2021; Janner et al., 2021; Emmons et al., 2022; David et al., 2023; Schmidhuber, 2019; Srivastava et al., 2019). This approach differs from the popular value-based methods (Kumar et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2022) by modeling the relationship between rewards and actions through supervised learning. Decision transformer (DT) (Chen et al., 2021) introduces the concept of desired future returns and improves performance by training the transformer architecture (Vaswani et al., 2017) as a return-conditioned policy. Based on DT, various advancements have been proposed for introducing value functions (Yamagata et al., 2023; Gao et al., 2024; Wang et al., 2024), finetuning models online (Zheng et al., 2022), adjusting the history length (Wu et al., 2023b), and improving the transformer architecture (Shang et al., 2022; Kim et al., 2024). However, these approaches do not explicitly address the alignment between the actual returns and the target return, which is a central focus of our work.
@@
Among the advancements based on DT, we delve deeper into studies that focus on modifying the model architecture similarly to our approach. Some efforts focus on refining the transformer architecture for offline RL. StARformer (Shang et al., 2022) introduces two transformer architectures, one aggregates information at each step, and the other aggregates information across the entire sequence. The image encoding process is improved by dividing the observation images into patches and feeding them into the transformer to enhance step information, similar to Vision Transformer (Dosovitskiy et al., 2021). Decision ConvFormer (Kim et al., 2024) replaces attention with convolution to capture the inherent local dependence pattern of MDP. These methods focus primarily on improving aspects of performance. To the best of our knowledge, this is the first work to explicitly align actual returns with target returns as the main objective in return-conditioned offline RL.
@@ 7 Discussion and Conclusion
In this paper, we proposed RADT, a novel decision-making model for aligning the actual return with the target return in offline RL. RADT splits the input sequence into return-to-go and state-action sequences, and reflects return-to-go in action generation by uniquely handling the return-to-go sequence. This unique handling includes two strategies that capture long-term dependencies and stepwise relationships within the return-to-go sequence. Experimental results demonstrated that RADT has superior aligning capabilities compared to existing DT-based models. One limitation of our method is a slight increase in computational cost compared to DT. This could potentially be improved by introducing a lightweight attention mechanism, such as Flash Attention (Dao et al., 2022), into our SeqRA. Another limitation is the difficulty in adapting to target returns that significantly exceed the maximum returns within the dataset. This issue could potentially be addressed by enhancing the training data returns through data augmentation techniques Li et al. (2024); Luo et al. (2025). We believe that RADT’s alignment capability will improve the usability of offline RL agents in a wide range of applications: creating diverse AI opponents in video games and educational tools, controlling heterogeneous agents in simulations, and efficiently generating human motions for animation or robotics. Investigating how RADT generalizes to multi-agent environments with diverse and adaptive opponents is an exciting avenue for future work.
Acknowledgments
This work was supported by JST, ACT-X Grant Number JPMJAX23CE, Japan. Kaito Ariu is supported by JSPS KAKENHI Grant No. 23K19986.